OpenStudy (anonymous):

If 8p = -2X2 - 100, which of the following could X possibly be?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[8p=-2x^2-100\]?

OpenStudy (anonymous):

Yes mam!

OpenStudy (misty1212):

so i guess your job is to solve for \(x\) right?

OpenStudy (anonymous):

Thanks for the help @misty1212 greatly appreciate it. And yes find X

OpenStudy (misty1212):

maybe start with \[8p+100=-2x^2\] then divide by \(-2\) and get \[-4p-50=x^2\]

OpenStudy (misty1212):

last step would be to take the square root and get \[x=\pm\sqrt{-4p-50}\]

OpenStudy (anonymous):

That's not a answer choice :9 This is so confusing.

OpenStudy (misty1212):

can you post a screen shot?

OpenStudy (misty1212):

or give the choices?

OpenStudy (anonymous):

x = -8p - 100 over 2

OpenStudy (anonymous):

How did you get that!?

OpenStudy (misty1212):

it it wrong we need a screen shot or sommat

OpenStudy (anonymous):

http://gyazo.com/0238f2b09b590917c2e11614eaf83973

OpenStudy (misty1212):

lol the idiots did not bother to divide by 2!!!

OpenStudy (misty1212):

go with C

OpenStudy (anonymous):

Thanks!!! Few more?

OpenStudy (misty1212):

sure i hope they get better

OpenStudy (anonymous):

Solve 3x2 + x + 10 = 0. Round solutions to the nearest hundredth. I got no real solutions but idk.

OpenStudy (misty1212):

no real solution looks good

OpenStudy (anonymous):

shes right

OpenStudy (anonymous):

Thanks @LesTwins

OpenStudy (anonymous):

Derek kicks a soccer ball off the ground and in the air, with an initial velocity of 31 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches?

OpenStudy (misty1212):

is it really 31, or could it be 32?

OpenStudy (anonymous):

idk -_- @satellite73

OpenStudy (misty1212):

i am asking if you made a typo is all it would be tons easier if it was "initial velocity of 32" instead of "31"

OpenStudy (anonymous):

31

OpenStudy (misty1212):

ok then if it is 31 it is 31 first coordinate of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\) which, in your case, is \[-\frac{31}{2\times( -16)}=\frac{31}{32}\]

OpenStudy (misty1212):

if it was initial velocity of 32 you would get a 1, which would be a lot easier to deal with, but whatever now you have to compute \[h(\frac{31}{32})\]

OpenStudy (misty1212):

i would use a calculator http://www.wolframalpha.com/input/?i=-16%2831%2F32%29%5E2%2B31%2831%2F32%29

OpenStudy (anonymous):

at max height H(t)=0 -32t+v=0 t=v/32 t=31/32 H(31/32)=-16(31/32)^2+31*31/32 = 15

OpenStudy (anonymous):

tahnks <3

OpenStudy (anonymous):

Last one A toy rocket was launched from the ground. The function f(x) = -16x2 + 128x shows the height of the rocket f(x), in feet, from the ground at time x seconds. What is the axis of symmetry of the graph of f(x), and what does it represent?

OpenStudy (anonymous):

@satellite73 pls